3) A rectangle with an area of 108 is divided into 3 equal squares. What is the perimeter of the rectangle?
Answer by xcentaur(357)

Visualize the problem,and you will be able to solve it much faster


let the length be x and width be y,of the rectangle.
then area=xy=108 unitsquare
no of squares=3
therefore area of each square=108/3=36 unitsquare
then side of each square is 6 units


the situation can be seen this way:

 --- --- ---
|   |   |   |
 --- --- ---

So we can see that the width of the rectangle is the side of the square=6 units and the length of the rectangles is 3 times the squares side,so the length is=3*6=18 units


So we have length=18,breadth=6
perimeter=2(x+y)=2(6+18)=2(24)=48 units


Hope this helps,
Prabhat